OpenStudy (anonymous):

QUESTION: −4(x + 10) − 6 = −3(x − 2) −40 −46 −52 52

OpenStudy (goten77):

solve for x i assume

OpenStudy (goten77):

\[-4(x + 10) - 6 = -3(x-2)\] 1st i distribute the -4 into its parenthesis and the -3 into its parenthesis and get \[-4x -40 -6 = -3x + 6\]

OpenStudy (goten77):

then college like terms \[-4x - 46 = -3x + 6\] but we also need just a single positive x on one side so we can add 4 x to each side getting us \[-46 = x + 6\] then add 6 to each side and we have x = -40